OpenStudy (anonymous):

Which of the following is equivalent to log 3x 5y ?

OpenStudy (anonymous):

Heres what I can tell you without the answer choices. If we have two values multiplied together in a log, then we can expand it into a sum of logs like this: \[\log (ab) = \log a + \log b\] so: \[\log 3x5y= \log 15xy = \log 15 + \log x + \log y\]

OpenStudy (anonymous):

alrighty, then what ? @vinnv226

OpenStudy (anonymous):

Thats as simplified as we can get. Although you could determine the value of log 15 if you'd like. If you'd like, you could post the answer choices and we could find the equivalent one.

OpenStudy (anonymous):

is the answer 4/3 ?

OpenStudy (anonymous):

We can't evaluate it unless we know what x and y are equal to. The log of 15 is equal to 1.176, if that helps you at all.

OpenStudy (anonymous):

no i think its 1.248

OpenStudy (anonymous):

nope, its 3/4

OpenStudy (anonymous):

Again, in order to reach an actual value, you need the values of x and y. Is there information you left out of the question?